OpenStudy (anonymous):

Can someone help me with, Select the inequality that corresponds to the given graph. (number line graph with closed circles on negative 1 and 3 with shading in between). (Without giving me the answer)!!!

OpenStudy (anonymous):

5x + 7 >= 2 and 10x <= 30 6x + 7 >= 25 or 2x - 8 < -4 4x + 10 <= 6 or 5x - 6 >= 14 37 <= 8x - 3 <= 61